### v1

# Sample from What You See: Visuomotor Policy Learning via Diffusion Bridge with Observation-Embedded Stochastic Differential Equation

Abstract
Imitation learning with diffusion models has advanced robotic control by capturing multi-modal action distributions. However, existing approaches typically treat observations as high-level conditioning inputs to the denoising network, rather than integrating them into the stochastic dynamics of the diffusion process itself. As a result, sampling must begin from random Gaussian noise, weakening the coupling between perception and control and often yielding suboptimal performance. We introduce BridgePolicy, a generative visuomotor policy that explicitly embeds observations within the stochastic differential equation via a diffusion-bridge formulation. By constructing an observation-informed trajectory, BridgePolicy enables sampling to start from a rich, informative prior rather than random noise, substantially improving precision and reliability in control. A key challenge is that classical diffusion bridges connect distributions with matched dimensionality, whereas robotic observations are heterogeneous and multi-modal and do not naturally align with the action space. To address this, we design a multi-modal fusion module and a semantic aligner that unify visual and state inputs and align observation and action representations, making the bridge applicable to heterogeneous robot data. Extensive experiments across 52 simulation tasks on three benchmarks and five real-world tasks demonstrate that BridgePolicy consistently outperforms state-of-the-art generative policies.
1 Introduction
Figure 1: Comparison of Diffusion Policy and our BridgePolicy. The observation modeling way of BridgePolicy allows its sampling of BridgePolicy can start from an rich and meaningful prior instead of the random noise in standard Diffusion Policy.
Figure 2: Overview pipelines. BridgePolicy explicitly embeds observations into the diffusion SDE trajectory via a diffusion-bridge formulation. The observation consists of robot states and point cloud. A multi-modal fusion module and a semantic aligner address the challenges of multi-modal distribution bridging and data shape mismatch, which enables effective exploitation of multi-modal observations. During the inference, BridgePolicy samples from the observation and iteratively transforms it into the action through fast sampling.
**footnotetext: Corresponding author.
Imitation learning (IL)  [osa2018algorithmic] is a widely adopted learning paradigm in robotic learning [li2024robust, shafiullah2023bringing, bin2020learning, ze2023gnfactor, seo2023deep, fu2024mobile], where a robot is provided with a set of expert demonstrations and learns to mimic the provided demonstrations to perform the tasks effectively. Recently, generative models such as diffusion model [ho2020denoising, score-based, classifier, classifier-free] and flow matching [lipman2024flow, flow_for_generative, flow_fast, SI] gains its prominence in IL owing to their capacity to fit multi-modal distributions and learn temporal dependency [chi2023diffusion, ze2023visual, zhang2025flowpolicy]. These methods share a similar principle: they perturb action trajectories into random noise via a forward process defined by a Stochastic or Ordinary Differential Equation (SDE/ODE) and then train a neural network conditioned on observations to reverse this process, iteratively transforming noise into executable actions.
With this paradigm, Diffusion Policy [chi2023diffusion] and 3D Diffusion Policy [ze2023visual], known as DP and DP3, respectively, employ an SDE-defined forward process and train the neural network conditioned on visual inputs and robot states to control the denoising process to sample actions from the random noise. FlowPolicy [zhang2025flowpolicy], in contrast, employs an ODE-defined forward and reverse process, which reduces the stochasticity during training and inference. Despite their success, current generative policies largely treat observations as high-level conditioning signals to the denoising network [chi2023diffusion, ze2023visual, zhang2025flowpolicy], rather than integrating them into the dynamics of the forward process. This underutilization forces sampling to begin from uninformative Gaussian noise, weakening the coupling between perception and control and often yielding suboptimal performance.
Diffusion bridge has demonstrated considerable success in the image generation and translation domain [yue2023image, li2023bbdm, luo2023image, de2021diffusion, zhou2023denoising], where it directly modifies the forward process such that the endpoint distribution naturally aligns with the desired conditioned distribution. As shown in Figure 1, with this mathematically exact formulation of modeling informative observations in the forward process rather than solely treating the observations as the external condition, the reverse process can naturally start from the observations, a more informative prior than the uninformative random noise, thereby improving the precision of the generated actions. Building upon these insights, we propose that diffusion bridge can also serve as a more effective framework for visuomotor policy learning. Specifically, rather than adopting the conventional diffusion model, we formulate policy learning as the problem of learning a diffusion bridge, where observations are explicitly modeled in to the diffusion SDE trajectory itself through the framework of diffusion bridge [zhu2025unidb, pan2025unidb] instead of merely treating them as conditions, and the reverse process would be able to sample actions starting from a more informative prior of the observations instead of the uninformative random Gaussian noise. This deeper integration allows the policy to more effectively exploit observations, leading to more precise and reliable control.
However, formulating policy learning as the diffusion bridge raises two key challenges. First, robotic observations and actions are inherently multi-modal and heterogeneous, violating the standard diffusion-bridge assumption that the connected endpoint distributions share the same dimensionality. Second, observations often include proprioceptive states, RGB-D vision, and language instructions, which do not admit a simple one-to-one mapping to the action space required by classical bridge formulations.
We introduce BridgePolicy, a generative visuomotor policy that directly samples actions from observation-informed priors rather than random noise. We construct a diffusion bridge that embeds observations within the diffusion SDE trajectory, enabling the model to fully exploit sensory inputs instead of using them solely as external conditions. To resolve the modality and shape mismatches, we design two components: (i) a multi-modal fusion module that consolidates heterogeneous inputs (e.g., vision, proprioception, language) into a unified observation representation, and (ii) a semantic aligner that maps this fused representation into an action-aligned latent space, ensuring compatible endpoints for the bridge. Together, these modules allow the diffusion bridge to leverage heterogeneous data sources for policy learning. Our contributions are summarized as follows:
  • We propose BridgePolicy, a novel generative visuomotor policy that explicitly embeds observations into the diffusion SDE trajectory via a diffusion-bridge formulation. This observation-informed SDE trajectory allows sampling to start from a meaningful prior rather than random noise, leading to more precise and reliable control.
  • To make diffusion bridges applicable to heterogeneous robotic data, we introduce a multi-modal fusion module and a semantic aligner that unify visual and state observations and align their representations with the action space. This resolves modality and shape mismatches, enabling effective use of multi-modal observations for policy learning.
  • Extensive experiments on 52 simulation tasks spanning three benchmarks and five real-world tasks show that BridgePolicy consistently outperforms prior generative policies, achieving state-of-the-art performance.
2 Related work
2.1 Generative Models in Robotic Learning
Recent years have witnessed the rise of diffusion models and flow matching as the state-of-the-art generative paradigms, largely driven by their remarkable success in the image generation and reinforcement learning domain [ho2020denoising, song2020denoising, rombach2022high, lipman2022flow, afforddp, QVPO, ding2025exploring, ding2025genpo], etc. Due to their capacity of multi-modal expressiveness, they are widely used in areas such as reinforcement learning, imitation learning, and motion planning [janner2022planning, reuss2023goal, chi2023diffusion, pearce2023imitating, sridhar2023memory, xian2023chaineddiffuser, prasad2024consistency, saha2024edmp]. In robotics, both diffusion models and flow matching have been utilized as policy learning frameworks, where action chunks are generated from a random noise under the condition of 2D or 3D observations through the SDE or ODE-defined generative processes, respectively [chi2023diffusion, ze20243d, zhang2025flowpolicy, hu2024adaflow]. In contrast, our method is pioneering in modeling the observations in the diffusion SDE trajectory via SOC. The sampling therefore starts from a more informative prior, allowing the policy to more effectively exploit observations and generate more precise and reliable behaviors.
2.2 Diffusion Bridges for Generative Modeling
Diffusion bridges enable the transition between two arbitrary distributions without the need to initiate the diffusion process from the random noise, which provides a more principled and mathematically exact generation paradigm for one-to-one distribution transformation. On the one hand, Diffusion Schrödinger Bridges [liu20232, shi2023diffusionschrodingerbridgematching, debortoli2023diffusionschrodingerbridgeapplications, somnath2024aligneddiffusionschrodingerbridges, chen2021likelihood, deng2024variational, wang2025implicit] aim to determine a stochastic process by solving an entropy-regularized optimal transport problem between two distributions. However, its high computational complexity, particularly pronounced in high-dimensional settings or constraints, poses significant challenges for direct optimization. Other works [zhou2023denoising, yue2023image] incorporate Doob’s $h$-transform into the forward SDE process, delivering remarkable performance on image restoration benchmarks. UniDB [zhu2025unidb] reformulates diffusion bridges through an SOC-based optimization problem, proving that Doob’s $h$-transform is a special case of SOC theory, thereby unifying and generalizing existing $h$-transform-based diffusion bridge methods. In this work, we leverage the ability of the diffusion bridge to connect heterogeneous distributions, specifically the multi-modal observation distribution and the action distribution of an expert policy, and harness this mechanism for policy learning.
3 Policy Learning via Diffusion Bridge
3.1 BridgePolicy
Existing policy learning frameworks [chi2023diffusion, ze20243d, zhang2025flowpolicy, hu2024adaflow] aim to learn a policy $\pi:\boldsymbol{O}\rightarrow\boldsymbol{A}$, which predict the actions $\boldsymbol{a}\in\boldsymbol{A}$ given the observations $\boldsymbol{o}\in\boldsymbol{O}$ which include 3D point cloud as the visual input and robot state and the actions are chunked into a short sequence of a trajectory finishing the task. Previous works such as DP3 [ze20243d] and FlowPolicy [zhang2025flowpolicy] formulate the policy as a conditional generative model, where actions are generated from random noises under the condition of observations $\boldsymbol{o}$ [chi2023diffusion, ze20243d, zhang2025flowpolicy]. They merely treated the informative observations as conditional signals to guide the neural network during denoising.
Forward Process of BridgePolicy. Motivated by the recent advances of Diffusion Bridge [zhou2023denoising, yue2023image, zhu2025unidb], where conditions can be explicitly modeled in the forward diffusion SDE trajectory, we directly integrate the observations into the diffusion process to generate a more controllable and precise robot action. Here, our BridgePolicy adopt a unified diffusion bridge framework [zhu2025unidb], formulated under the Stochastic Optimal Control (SOC) theory, as the core learning paradigm. To be consistent with the notations of previous diffusion for visuomotor policy learning [chi2023diffusion], we denote $\boldsymbol{a}_{t}$ as the interpolated robot action state at time $t$. Assuming that the action $\boldsymbol{a}$ and the observation $\boldsymbol{o}$ share the same dimension, we define the endpoints of our forward diffusion bridge process (2) as follows:
  • The initial state $\boldsymbol{a}_{0}$: we set the initial state to be the action, i.e., $\boldsymbol{a}_{0}=\boldsymbol{a}$.
  • The terminal state $\boldsymbol{a}_{T}$: we set the terminal state to be the observation, i.e., $\boldsymbol{a}_{T}=\boldsymbol{o}$. It is a pivotal departure from standard diffusion-based or flow-based policy where $\boldsymbol{a}_{T}\sim\mathcal{N}(0,I)$ is typically random Gaussian noise.
According to UniDB [zhu2025unidb], we can construct the forward process of diffusion bridge as the following SOC optimization problem:
Algorithm 1 BridgePolicy Training
Input: Actions $\boldsymbol{a}$, Observations $\boldsymbol{o}=\{\boldsymbol{o}_{s},\boldsymbol{o}_{pc}\}$, loss weight $\alpha$, and gradient descent step size $\eta$.
repeat
  $\boldsymbol{z}_{s}=\mathbf{MLP}_{\phi_{1}}(\boldsymbol{o}_{s})$, $\boldsymbol{z}_{pc}=\mathbf{MLP}_{\phi_{2}}(\boldsymbol{o}_{pc})$
  $\boldsymbol{z}_{obs}=\mathbf{softmax}({\boldsymbol{z}_{s}\cdot\boldsymbol{z}_{s}^{T}}/{\sqrt{d_{s}}})\boldsymbol{z}_{pc}$
  $\boldsymbol{a}_{0}=\boldsymbol{a},$ $\boldsymbol{a}_{T}=\boldsymbol{z}_{obs}$
  $t\sim\text{Uniform}(\{1,...,T\})$
  Compute $\mathcal{L}=\mathcal{L}_{DB}+\alpha\mathcal{L}_{align}$
  Update $(\theta,\phi_{1},\phi_{2})\leftarrow(\theta,\phi_{1},\phi_{2})-\eta\nabla\mathcal{L}$
until Converged
$\begin{gathered}\min_{\mathbf{u}_{t,\gamma}}\ \mathbb{E}\left[\int_{0}^{T}\frac{1}{2}\left\|\mathbf{u}_{t,\gamma}\right\|_{2}^{2}dt+\frac{\gamma}{2}\left\|\boldsymbol{a}_{T}^{u}-\boldsymbol{a}_{T}\right\|_{2}^{2}\right]\\ \text{s.t.}\ \mathrm{d}\boldsymbol{a}_{t}=\Big[\theta_{t}(\boldsymbol{a}_{T}-\boldsymbol{a}_{t})+g_{t}\mathbf{u}_{t,\gamma}\Big]\mathrm{d}t+g_{t}\mathrm{d}\boldsymbol{w}_{t},\ \boldsymbol{a}_{0}^{u}=\boldsymbol{a}_{0},\end{gathered}$ (1)
where $\boldsymbol{a}_{0}^{u}$ and $\boldsymbol{a}_{T}^{u}$ are the endpoints under control to be distinguished with the pre-defined endpoints, $\theta_{t}$ and $g_{t}$ are scalar-valued functions with the relation $g_{t}^{2}=2\lambda^{2}\theta_{t}$ and the steady variance level $\lambda^{2}$ is a given constant, $\boldsymbol{w}_{t}$ denotes the Wiener process, $\left\|\mathbf{u}_{t,\gamma}\right\|_{2}^{2}$ is the trajectory cost, and $\frac{\gamma}{2}\left\|\boldsymbol{a}_{T}^{u}-\boldsymbol{a}_{T}\right\|_{2}^{2}$ is the terminal cost with its penalty coefficient $\gamma$. The SOC problem (1) targets to design the optimal controller $\mathbf{u}^{*}_{t,\gamma}$ to drive the dynamic system from $\boldsymbol{a}_{0}$ to $\boldsymbol{a}_{T}$ with minimum cost. UniDB solved the problem (1) and obtained a closed-form optimally controlled forward SDE to transform $\boldsymbol{a}_{0}$ into $\boldsymbol{a}_{T}$ as:
$\begin{gathered}\mathrm{d}\boldsymbol{a}_{t}=\left(\theta_{t}+\frac{g^{2}_{t}e^{-2\bar{\theta}_{t:T}}}{\gamma^{-1}+\bar{\sigma}^{2}_{t:T}}\right)(\boldsymbol{a}_{T}-\boldsymbol{a}_{t})\mathrm{d}t+g_{t}\mathrm{d}\boldsymbol{w}_{t},\end{gathered}$ (2)
where $\bar{\theta}_{s:t}=\int_{s}^{t}\theta_{z}dz$, $\bar{\theta}_{t}=\int_{0}^{t}\theta_{z}dz$, $\bar{\sigma}^{2}_{s:t}=\lambda^{2}(1-e^{-2\bar{\theta}_{s:t}})$, and $\bar{\sigma}^{2}_{t}=\lambda^{2}(1-e^{-2\bar{\theta}_{t}})$. Through the diffusion bridge paradigm, we successfully construct the mapping from actions $\boldsymbol{a}_{0}=\boldsymbol{a}$ to observations $\boldsymbol{a}_{T}=\boldsymbol{o}$.
Training Objective of BridgePolicy. In order to learn such a bridge to reverse the process, we adopt the following reconstruction training objective to train a data prediction neural network $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$ that directly maps any $\boldsymbol{a}_{t}$ to the clean action $\boldsymbol{a}_{0}$:
$\mathcal{L}_{DB}=\mathbb{E}_{t,\boldsymbol{a},\boldsymbol{a}_{T},\boldsymbol{a}_{t}}\left[\|\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)-\boldsymbol{a}\|\right],$ (3)
where we follow the $l_{1}$ norm in UniDB [zhu2025unidb] regarding the order of the norm, and provide the ablation study in the Supplementary Materials. Consequently, BridgePolicy naturally integrates the informative observations $\boldsymbol{a}_{T}=\boldsymbol{o}$ both in the SDE trajectories and the neural networks and aims to learn a stochastic trajectory that transforms the observation $\boldsymbol{o}$ into a plausible action $\boldsymbol{a}$.
Algorithm 2 BridgePolicy Inference
Input: Observations $\boldsymbol{o}=\{\boldsymbol{o}_{s},\boldsymbol{o}_{pc}\}$, pre-trained data prediction model $\boldsymbol{a}_{\theta}$, and $M+1$ time steps $\left\{t_{i}\right\}_{i=0}^{M}$ decreasing from $t_{0}=T$ to $t_{M}=0$.
$\boldsymbol{z}_{s}=\mathbf{MLP}_{\phi_{1}}(\boldsymbol{o}_{s})$, $\boldsymbol{z}_{pc}=\mathbf{MLP}_{\phi_{2}}(\boldsymbol{o}_{pc})$
 Initialize $\boldsymbol{a}_{T}=\boldsymbol{z}_{obs}=\mathbf{softmax}({\boldsymbol{z}_{s}\cdot\boldsymbol{z}_{s}^{T}}/{\sqrt{d_{s}}})\boldsymbol{z}_{pc}$
for $i=1$ to $M$ do
  Sample $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ if $i<M$, else $\boldsymbol{\epsilon}=0$
  $\boldsymbol{a}_{t_{i}}\leftarrow\text{Update}(\boldsymbol{a}_{t_{i-1}},\boldsymbol{a}_{T},\boldsymbol{a}_{\theta},\boldsymbol{\epsilon},t_{i-1})$ from (4)
end for
Return Actions $\tilde{\boldsymbol{a}}$
Fast Sampling Algorithm. After we obtain the optimal model $\boldsymbol{a}^{*}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$, the decision making module of our BridgePolicy can be implemented as an iterative generation procedure starting from the given observation $\boldsymbol{a}_{T}=\boldsymbol{o}$. As for specially designed solvers, UniDB++ [pan2025unidb] provides a training-free acceleration algorithm with the data prediction $\boldsymbol{a}_{\theta}(\boldsymbol{a}_{t},\boldsymbol{a}_{T},t)$ which directly estimates the fixed and smooth target $\boldsymbol{a}_{0}$. With some notations $\rho_{t}=e^{\bar{\theta}_{t}}(1-e^{-2\bar{\theta}_{t}})$, $\kappa_{t,\gamma}=e^{\bar{\theta}_{t:T}}((\gamma\lambda^{2})^{-1}+1-e^{-2\bar{\theta}_{t:T}})$, $c_{1}=(\gamma\lambda^{2})^{-1}e^{2\bar{\theta}_{T}}$, $c_{2}=e^{2\bar{\theta}_{T}}-1$, $D=2c_{1}c_{2}/(c_{1}+c_{2})^{3}$, $E=c_{2}^{2}/(c_{1}+c_{2})^{2}$, and $F=c_{1}^{2}/(c_{1}+c_{2})^{2}$ in UniDB++, the closed-form updating rule from time step $t$ to $s$ is formed as
$\begin{gathered}\begin{aligned} \boldsymbol{a}_{t}=&\frac{\kappa_{t,\gamma}\kappa_{s}\rho_{t}}{\kappa_{s,\gamma}\kappa_{t}\rho_{s}}\boldsymbol{a}_{s}+(1-\frac{\kappa_{t,\gamma}\kappa_{s}\rho_{t}}{\kappa_{s,\gamma}\kappa_{t}\rho_{s}}+\frac{\kappa_{t,\gamma}\rho_{t}\kappa_{s}}{\kappa_{0,\gamma}\kappa_{t}\rho_{s}}-\frac{\kappa_{t,\gamma}}{\kappa_{0,\gamma}})\boldsymbol{a}_{T}\\ &+\left(\frac{\kappa_{t,\gamma}}{\kappa_{0,\gamma}}-\frac{\kappa_{t,\gamma}\rho_{t}\kappa_{s}}{\kappa_{0,\gamma}\kappa_{t}\rho_{s}}\right)\boldsymbol{a}_{\theta}(\boldsymbol{a}_{s},\boldsymbol{a}_{T},s)+\delta^{d}_{s:t}\boldsymbol{\epsilon},\end{aligned}\\ \begin{aligned} (\delta^{d}_{s:t})^{2}&=\frac{\lambda^{2}\kappa^{2}_{t,\gamma}\rho^{2}_{t}}{\kappa^{2}_{t}}\Bigg[E\frac{e^{2\bar{\theta}_{s}}-e^{2\bar{\theta}_{t}}}{(e^{2\bar{\theta}_{t}}-1)(e^{2\bar{\theta}_{s}}-1)}\\ &+D\log\frac{\kappa_{s,\gamma}\rho_{t}}{\kappa_{t,\gamma}\rho_{s}}-F(\frac{e^{-\bar{\theta}_{T}-\bar{\theta}_{t}}}{\kappa_{t,\gamma}}-\frac{e^{-\bar{\theta}_{T}-\bar{\theta}_{s}}}{\kappa_{s,\gamma}})\Bigg],\end{aligned}\end{gathered}$ (4)
where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ is the Guassian noise. Although the updating rule seems very complex on the coefficients of $\boldsymbol{a}_{s}$, $\boldsymbol{a}_{T}$, $\boldsymbol{a}_{\theta}$, and $\boldsymbol{\epsilon}$, actually these coefficients can be computed explicitly without incurring significant computational overhead, because they have closed forms and depend only on the specific time steps.
3.2 Modality Fusion and Alignment
While conceptually appealing under a uniform dimensionality assumption, directly applying the diffusion bridge paradigm remains infeasible, as such an assumption rarely holds in practice. This method still faces challenges of multi-modal distribution bridging and data shape mismatch, which prevent the diffusion bridge from functioning effectively as a policy learning method. To address these limitations, we propose a novel framework that effectively bridges heterogeneous modalities through a carefully designed encoding and fusion pipeline.
Encoding the Robot State and Point Cloud. Our encoding pipeline begins by converting depth images into 3D point clouds, chosen for the efficiency. We then downsample them (512 or 1024 points in simulation; 2048 in real-world) using farthest point sampling [qi2017pointnet]. Finally, a lightweight MLP with LayerNorm and an MLP state encoder map the point cloud and robot state into a shared latent representation for fusion, which can be formulated as:
$\boldsymbol{z}_{s}=\mathbf{MLP}_{\phi_{1}}(\boldsymbol{o}_{s}),\quad\boldsymbol{z}_{pc}=\mathbf{MLP}_{\phi_{2}}(\boldsymbol{o}_{pc}),$ (5)
where $\phi_{1}$ and $\phi_{2}$ are the MLP network parameters, $\boldsymbol{o}_{s}$ is the robot state, and $\boldsymbol{o}_{pc}$ represents the point cloud.
Figure 3: Examples of the simulation tasks.
Multi-Modality Fusion. To enable the diffusion bridge to map between heterogeneous modalities effectively, we introduce the multi-modality fusion module that integrates the 3D point cloud and robot state into a unified representation. This allows the policy to process diverse inputs coherently within our framework. Specifically, we employ the Cross-Attention mechanism [rombach2022high] to fuse the modalities:
$\boldsymbol{a}_{T}:=\boldsymbol{z}_{obs}=\mathbf{softmax}(\frac{\boldsymbol{z}_{s}\boldsymbol{z}_{s}^{T}}{\sqrt{d_{s}}})\boldsymbol{z}_{pc},$ (6)
where $\boldsymbol{z}_{s}$ represents the feature vector of the robot state, $d_{s}$ is the dimension of $\boldsymbol{z}_{s}$, and $\boldsymbol{z}_{pc}$ denotes the feature vector of the visual input. Here, $\boldsymbol{a}_{T}=\boldsymbol{z}_{obs}$ is the unified observation feature vector with the same shape of the action chunk, representing the unified observation representation.
Figure 4: Examples of the real-world tasks.
Modality Alignment. Although multi-modality fusion aligns the observation and action distributions in shape, significant distributional differences still remain and the learned observation representation lacks sampling capability. To address this, we propose using the contrastive learning loss to train the aligner. The contrastive loss aligns the semantic proximity of the observation and action distributions. Specifically, we adopt CLIP loss [radford2021learning] as the contrastive learning loss:
Table 1: Main Simulation Results. Quantitative comparison on success rates among the baselines and our BridgePolicy. We evaluate 50 tasks across 3 benchmarks and report the average success rates of each benchmarks. For MetaWorld, we group the tasks based on their difficulty levels and report the average success rates.
[…]
Easy Medium Hard Very Hard
DP [chi2023diffusion] 0.79±0.27 0.31±0.24 0.10±0.12 0.26±0.25 0.45±0.08 0.31±0.17 0.37
DP3 [ze20243d] 0.87±0.20 0.61±0.29 0.40±0.37 0.51±0.33 0.57±0.08 0.68±0.06 0.60
Simple DP3 [ze20243d] 0.86±0.12 0.59±0.22 0.38±0.26 0.47±0.28 0.48±0.07 0.68±0.05 0.58
FlowPolicy [zhang2025flowpolicy] 0.86±0.16 0.67±0.21 0.59±0.30 0.76±0.26 0.54±0.09 0.70±0.12 0.68
[…]
Figure 5: Real-robot comparative visualization of DP3, FlowPolicy, and BridgePolicy at four critical waypoints for Oven-Opening task.
$\begin{gathered}\mathcal{L}_{align}=\mathcal{L}_{clip}(\boldsymbol{a},\boldsymbol{z}_{obs})+\mathcal{L}_{clip}(\boldsymbol{z}_{obs},\boldsymbol{a}),\\ \mathcal{L}_{clip}(\boldsymbol{x},\boldsymbol{y})=-\frac{1}{n}\sum_{j=1}^{n}\log\frac{\exp(x_{j}^{\top}y_{j}/\tau)}{\sum_{i=1}^{n}\exp(x_{i}^{\top}y_{j}/\tau)},\end{gathered}$ (7)
where $n$ is the batch size and $\tau$ is the temperature parameter. The entire model involves the diffusion bridge and the modality fusion module and therefore, the overall training objective integrates the diffusion bridge loss (3) with the alignment loss (7):
$\mathcal{L}=\mathcal{L}_{DB}+\alpha\mathcal{L}_{align},$ (8)
where $\alpha$ is a positive weight. We provide the pseudo-code Algorithm 1 and Algorithm 2 for the training and inference process of our BridgePolicy. During the training phase, the observations are firstly fused and reshaped by the modality fusion module and aligner, and then alignment loss and diffusion bridge loss are computed by corresponding equations. After training, the observations are fused into the latent vector set as the starting point of the sampling process. This latent vector is subsequently updated iteratively by the designed solver (4), ultimately generating the actions.
4 Experiment
4.1 Simulation Experiment
Table 2: Main Real-world Results. Quantitative comparison of success rates on real-world tasks among the baselines and BridgePolicy.
Oven-Closing Oven-Opening Pick Place Pour Unplug Average
Simple DP3 0.8 0.6 0.6 0.6 0.7 0.66
DP3 0.9 0.9 0.7 0.6 0.7 0.76
FlowPolicy 1.0 0.7 0.5 0.1 0.5 0.56
[…]
Simulation Benchmark. In simulation tasks, we evaluate our BridgePolicy on 52 tasks across three benchmarks including Adroit [rajeswaran2017learning], DexArt [bao2023dexart], and MetaWorld [yu2020meta]. Adroit and DexArt focus on dexterous hand manipulation with varying complexity, while MetaWorld offers a broad spectrum of robotic arm tasks across different difficulty levels. Examples of tasks on these three benchmarks are illustrated in Figure 3.
Expert Demonstration Collection. We collect expert data with script policy in MetaWorld and adopt the VRL3 [wang2022vrl3] and PPO [schulman2017proximal], two reinforcement learning (RL) algorithms, to collect expert data for Adroit and DexArt, respectively. We collect 10 episodes for tasks of the Adroit and the MetaWorld benchmarks and 100 episodes for the DexArt benchmark to train the policies.
Baselines. For comparison, we select state-of-the-art 2D-based methods DP [chi2023diffusion] which takes the 2D image as its visual input and 3D conditional diffusion or flow based approaches including DP3 [ze20243d], simple DP3 (its lightweight version), and FlowPolicy [zhang2025flowpolicy], which generates the action from random Gaussian noise and leverage 3D point cloud as the visual condition in the neural network as baselines. We set the number of function evaluation (NFE) to 10 for all baselines except FlowPolicy. Based on the settings in the original FlowPolicy paper, we set its NFE to 1 instead of 10 to avoid the error accumulation issue that occurs in multi-step sampling of consistency flow matching [sabour2025align]. For further details of the NFE setting, please refer to the Supplementary Materials.
Evaluation and Implementation. We run 3 random seeds for each experiment, followed by DP3 [ze20243d] and FlowPolicy [zhang2025flowpolicy]. For each random seed, we train the models for 3000 epochs and evaluate the success rate on 20 episodes every 200 training epochs. We pick the five highest success rates per seed and report the mean success rate of the 3 random seeds. For fair comparison, we keep the model architecture consistent with the baselines, the amount of parameters comparable, and the shared hyperparameters, such as learning rates, weight decays of the optimizer, batch sizes, and training epochs. For each task, BridgePolicy is trained on a single NVIDIA RTX 4090 GPU. For further details of the implementation of BridgePolicy and other baselines, please refer to the Supplementary Materials.
Performance in Simulator. For Adroit and DexArt, we report the average success rates directly. Particularly for MetaWorld tasks, it is further categorized into four difficulty levels, with average rates reported for each. The quantitative results are reported in Table 1. Our proposed BridgePolicy demonstrates a consistent performance across all evaluated benchmarks, achieving the highest average success rates. All baseline methods show considerably lower success rates, particularly in more challenging settings (Hard and Very Hard in MetaWorld), which also showcases the robustness of our method in tasks with varying complexity. These results validate the effectiveness of BridgePolicy in leveraging the diffusion bridge formulation for improved policy learning in diverse simulation environments. For success rates of individual tasks, please refer to the Supplementary Materials.
Figure 6: Ablation on the number of demonstrations. BridgePolicy shows its superiority with less training demonstrations.
4.2 Real-world Experiment
Table 3: Quantitative evaluation results (Success Rate) on different modality fusion methods (MetaWorld is denoted as MW in table).
Task Modality Fusion Method
Concatenation Cross-Attention
Adroit Pen 0.78 0.81
Adroit Door 0.59 0.665
MW Handle-Pull 0.55 0.63
Experiment Setup and Evaluation. We evaluate the BridgePolicy on 5 real-world tasks using the Franka Emika Panda robot. The point cloud is acquired using the ZED-2i camera. The real-world tasks are shown in Figure 4. We evaluate the policies on 10 episodes of each task. The implementation remains the same as that in the experiments in simulation. We only increase the number of points of the point cloud to 2048 for a more dense representation of the real-world scenario. For experiment environment setup details of the real-world experiments, please refer to the Supplementary Materials.
Expert Demonstration Collection. The expert demonstrations are collected by the GELLO human teleopertation system [wu2024gello], manipulated by an experienced graduate. For each task, we collect 50 episodes to train the policy. Similarly, in the simulator and for a fair comparison, the expert demonstrations provided to all baseline methods and our BridgePolicy are identical. Please refer to the Supplementary Materials for setup details of the real-world data collection.
Quantitative Comparison on Real-World Task. The success rates of real-world tasks are shown in Table 2. Under the identical constraint of training with the same dataset of 50 episodes, our BridgePolicy yields the highest success rate on all the tasks, with an average of 0.9, underscoring its effectiveness. Meanwhile, FlowPolicy exhibits a significant performance drop in the real world on tasks requiring location generalization, such as Pour and Pick-and-Place. DP3 is in the second place with a success rate of 0.78, and Simple DP3 only achieves a success rate of 0.65, possibly due to the small number of parameters.
Qualitive Comparison on Real-World Task. As depicted in Figure 5 and  7, we select three examples of the real-world task Oven-Opening and Unplug generated by DP3, FlowPolicy, and BridgePolicy, respectively. From left to right, the sequence presents the key frames of the Franka arm finishing the task. The Oven-Opening task consists of 2 key stages: 1) grabbing the door handle of the oven and opening it halfway; 2) moving the arm upon the half-opened door and pushing it down until it is fully opened. In the very case of Oven-Opening, DP3 fails to finish the task while FlowPolicy and BridgePolicy both finish the task in the end. BridgePolicy exactly finishes the 2 stages and fully opens the oven, while FlowPolicy fails to half-open the door in the first stage, and it presses down against the door handle directly, leaving the door not fully opened. We also observe similar results in the Unplug task, shown in Figure 7. DP3 and FlowPolicy are able to extend their grippers towards the target location, but they did not precisely clamp the socket, thus failing to unplug successfully. BridgePolicy, on the other hand, successfully clamped the plug and pulled it out. Videos of the real-world experiments can be found in the Supplementary Materials.
Table 4: Quantitative evaluation results (Success Rate) on different Adroit tasks with different loss weights $\alpha$.
Task Loss weights $\alpha$
0.3 0.6 1.0
Adroit Door 0.81 0.88 0.81
Adroit Pen 0.665 0.615 0.61
MetaWorld Box-Close 0.74 0.75 0.76
MetaWorld Coffee-Pull 0.95 0.86 0.84
MetaWorld Dial-Turn 0.91 0.87 0.85
MetaWorld Handle-Pull 0.63 0.53 0.49
Avg Success Rate 0.78 0.75 0.72
4.3 Ablation Study
In this section, we analyze the performance of BridgePolicy under different hyperparameter settings. Additional ablation studies are provided in the Supplementary Materials.
Sample Efficiency. The success rate of the agent in accomplishing tasks largely depends on the number of expert demonstrations. Here, we evaluate how the number of expert demonstrations would affect the success rate of the agent finishing the tasks. As shown in Figure 6, we select four tasks for this ablation study: the MetaWorld Hand-Pull-Side, Pick Place, Stick Pull and the Adroit Pen task. For all four tasks, increasing the number of expert demonstrations generally improves the agent’s success rate. BridgePolicy shows the most stable and robust performance across all demonstration counts among the evaluated tasks, especially with fewer training samples. For more sample efficiency ablations, please refer to the Supplementary Materials.
Parameter Sensitivity. We demonstrate how the key parameters of the BridgePolicy would influence its performance. Specifically, we evaluate different weights $\alpha$ in the overall training loss (8) and the results are shown in Table 4. The results suggest that the distribution aligner can be trained better under a relatively small weight in policy learning. For more ablation study results on the hyperparameters, please refer to the Supplementary Materials.
Modality Fusion. We additionally explore the different ways of fusing the different modalities of the observations. Modality fusion plays a crucial role in resolving the conflict between the multi-modal nature of the observations in policy learning and the one-to-one distribution mapping of the diffusion bridge. We explore the impact of two modality fusion methods, Cross-Attention [lin2022cat] and simple concatenation, on the performance. The results are shown Table 3. When switching from Cross-Attention to simple concatenation, the performance consistently drops on all three evaluated tasks, which demonstrates the effectiveness of using Cross-Attention.
Figure 7: Real-robot comparative visualization of DP3, FlowPolicy, and BridgePolicy at four critical waypoints for Unplug task. BridgePolicy successfully grabbed the plug, while DP3 and FlowPolicy can only move to the approximate location.
Figure 8: Real-robot comparative visualization of DP3, FlowPolicy, and BridgePolicy at four critical five points for Pick and Place task.
Figure 9: Real-robot comparative visualization of DP3, FlowPolicy, and BridgePolicy at five critical points for Pick and Place task.
5 Conclusion
In this work, we introduce BridgePolicy, a novel generative policy learning paradigm that directly embeds the observations in the diffusion SDE trajectory via a diffusion bridge formulation rather than merely treating them as high-level conditions in the denoising network like conventional generative policies. This observation-informed SDE enables the sampling to start from a rich and meaningful prior, allowing the policy to more effectively exploit observations, leading to more precise and reliable control. Extensive experiments on simulation and real-world tasks demonstrate its superiority over the current generative policy. The design choices for the distribution aligner and the way of modality fusion, while effective in our experiments, represent one point in a broader design space. We see great potential in further refining these components, and future work will focus on exploring alternative architectures for even more robust distribution alignment and seamless modality integration. In the end, we hope our method helps pave the way towards further research and applications of the diffusion bridge and visuomotor policy learning.
Acknowledgment
This work was supported by the National Natural Science Foundation of China (62303319, 62406195), Shanghai Local College Capacity Building Program (23010503100), ShanghaiTech AI4S Initiative SHTAI4S202404, HPC Platform of ShanghaiTech University, and MoE Key Laboratory of Intelligent Perception and Human-Machine Collaboration (ShanghaiTech University), Shanghai Engineering Research Center of Intelligent Vision and Imaging. This work was also supported in part by computational resources provided by Fcloud CO., LTD.
Sample from What You See: Visuomotor Policy Learning via Diffusion Bridge with Observation-Embedded Stochastic Differential Equation
Supplementary Materials
Appendix A Overview
We provide additional information, results, and visualization to supplement the main paper. The supplementary materials are organized as follows:
  • Information of experimental setup and real-world data collection;
  • Training and inference details;
  • Implementation details of BridgePolicy;
  • Additional and expanded experimental results.
Appendix B Information of experimental setup and real-world data collection
B.1 Real-World Experiment Setup
The real-world experiment was conducted on a Franka arm. We equipped it with two types of grippers, one from Robotiq and the other from FastUMI. The specific scenario setup is shown in Figure 12. The FastUMI gripper is used to finish the Pour, Oven-Opening, Oven-closing and Pick-and-Place tasks and the Robotiq gripper is used to finish the Unplug task.
Figure 10: The GELLO human teleoperate system for real-world data collection.
Figure 11: The visualization of point cloud of the key frames of the Unplug task.
B.2 Data Collection
We use the GELLO human teleoperate system [wu2024gello] to collect the data for real-world experiments. The GELLO controller device and the GELLO data collection system we use is illustrated in Figure 10. We perform hand-eye calibration in the real-world data collection. The calibration process enabled us to precisely determine the transformation relation from the camera coordinate system to the robot-base coordinate system. The observations including 3D point clouds and robot states and actions are recorded and consist the expert demonstration dataset. We use 3D point clouds as the robot’s visual input, which is directly acquired by the ZED-2i camera, a binocular camera can generate the point cloud of the scene. An example of the point cloud acquired by the camera is shown in Figure 11. A single point cloud output from the ZED-2i camera contains 252,672 points. The point cloud input to the policies including BridePolicy and baselines was further downsampled to 2048 points using the Farthest Point Sampling (FPS) algorithm and done by the fpsample library for further acceleration to support the real-time inference.
Figure 12: The experimental setup for the unplug task with Robotiq and FastUMI gripper.
Table 5: Evaluation of FlowPolicy using 10 step and 1 step when sampling
Flow (1 step) Flow (10 step)
Adroit Hammer 1.00 0.91
Adroit Door 0.53 0.50
DexArt Faucet 0.38 0.37
DexArt Bucket 0.32 0.26
MetaWorld Hammer 0.98 0.94
MetaWorld Coffee-Pull 0.96 0.89
MetaWorld Soccer 0.33 0.24
Average 0.64 0.59
Appendix C Model Implementation of BridgePolicy and Baselines
C.1 Implementation of The Denoising Network
Following the implementation of the denoising model in DP3 [ze20243d] and FlowPolicy [zhang2025flowpolicy], we adopt the U-net architecture as the backbone of the denoising model in BridgePolicy and the baselines. The numbers of parameters in the denoising model in BridgePolicy and the baselines are the same which is approximately 255M. The BridgePolicy possess additional modules, the semantic aligner and the modality fusion module, which only have about less than 1M parameters.
Appendix D Training and Inference Details of BridgePolicy and Baselines
D.1 Training
For the configurations such as the horizon length, the number of observation steps and the predicted action steps, we set them to the default setting in their publicly available code base and the corresponding configurations of BridgePolicy is aligned with DP3. Specifically, we list the hyperparameters used in BridgePolicy in Table 6 (Real-World).
Table 6: Hyperparameter setting in BridgePolicy in real-world experiments.
Hyperparameter Value
Num epochs 4000
Batch Size 256
Horizon 8
Observation Steps 2
Action Steps 4
Num points 2048
Num train timesteps 100
Num inference steps 10
Learning Rate (LR) 1.0e-4
Aligner weight $\alpha$ 1.0
Weight decay 1.0e-6
Table 7: Ablation Study: Success Rates of Different $\alpha$ on MetaWorld
$\alpha$ MetaWorld Avg
basketball box-close coffee-pull coffee-push dial-turn hammer handle-pull pick-place-wall
0.1 1.00 0.68 0.95 0.94 0.92 1.00 0.59 0.98 0.88
0.3 1.00 0.72 0.95 0.93 0.91 0.97 0.53 0.94 0.87
0.5 1.00 0.75 0.88 0.93 0.87 0.97 0.56 0.93 0.86
0.6 1.00 0.75 0.96 0.90 0.93 0.97 0.54 0.91 0.87
0.8 1.00 0.81 0.96 0.83 0.85 0.98 0.53 0.93 0.86
1.0 1.00 0.66 0.92 0.92 0.85 0.99 0.57 0.93 0.85
1.2 1.00 0.86 0.91 0.91 0.78 0.96 0.56 0.93 0.86
1.5 1.00 0.75 0.92 0.93 0.78 0.96 0.56 0.88 0.85
D.2 Inference
Number of Sampling Steps. We set the number of sampling steps to be 10 for BridgePolicy and baselines except the number of sampling steps of FlowPolicy is set to 1 according to the original paper [zhang2025flowpolicy]. We also conduct additional experiments to explore how the performance would change if we increase the number of sampling steps and the results are shown in Table 5. These results show that the performance degrades, possibly due to the error accumulation issue in consistency flow matching [sabour2025align].
Appendix E Additional Experimental Results
We provide additional experimental results in this section. Despite the more ablation results and success rates of individual tasks, we also provide additional visualization of the real-world tasks shown in Figure 9 and 8.
E.1 More Ablation Results on $\alpha$ parameter
We provide more ablation results on the $\alpha$ parameter shown in Table 7 and 8. For most of the evaluated tasks, a smaller $\alpha$ less than $1$ is better for training the policy.
Table 8: Ablation Study: Success Rates of Different $\alpha$ on Adroit and DexArt.
$\alpha$ Adroit Adroit Avg DexArt DexArt Avg
hammer door pen laptop faucet toilet bucket
0.1 1.00 0.84 0.60 0.81 0.87 0.41 0.73 0.35 0.59
0.3 1.00 0.79 0.67 0.82 0.88 0.43 0.75 0.33 0.60
0.5 1.00 0.85 0.56 0.80 0.83 0.45 0.71 0.33 0.58
0.6 1.00 0.84 0.52 0.78 0.88 0.45 0.76 0.30 0.60
0.8 1.00 0.81 0.62 0.81 0.87 0.42 0.74 0.37 0.60
1.0 1.00 0.88 0.58 0.82 0.89 0.42 0.73 0.31 0.59
1.2 1.00 0.80 0.61 0.80 0.85 0.44 0.73 0.30 0.58
1.5 1.00 0.81 0.60 0.80 0.85 0.42 0.73 0.31 0.57
E.2 Parameter Sensitivity on $\gamma$ and $\lambda$
We demonstrate how the key parameters of the BridgePolicy would influence its performance. Specifically, we evaluate the combinations of two hyperparameters (the steady variance level $\lambda^{2}$ (over 255) and the penalty coefficient $\gamma$) in the UniDB [zhu2025unidb] framework and the results are shown in Table 9. Results in Table 9 demonstrates that $\lambda=50$ and $\gamma=10^{7}$ are the relatively better hyperparameter combination of the UniDB framework in policy learning.
Table 9: Quantitative evaluation results on different Adroit tasks with different steady variance levels $\lambda^{2}$ and penalty coefficients $\gamma$.
Task Hyperparameters ($\lambda,\gamma$)
(30, $10^{7}$) (50, $10^{7}$) (70, $10^{7}$) (30, $10^{5}$) (50, $10^{5}$) (70, $10^{5}$)
Adroit Hammer 0.9 1.0 1.0 0.85 1.0 1.0
Adroit Door 0.75 0.85 0.85 0.65 0.85 0.825
Adroit Pen 0.65 0.75 0.65 0.65 0.65 0.6
Avg Success Rate 0.77 0.87 0.83 0.72 0.83 0.81
E.3 Success Rates of Individual Tasks
We list the success rate of individual task of each benchmark in Table 10. All results are averaged over three random seeds.
Table 10: Success Rates of Individual Task on Adroit and DexArt benchmarks.
Adroit DexArt
[…]
DP [chi2023diffusion] 0.45±0.05 0.35±0.04 0.15±0.04 0.69±0.08 0.23±0.09 0.63±0.08 0.23±0.08
DP3 [ze20243d] 1.00±0.00 0.65±0.05 0.40±0.05 0.86±0.08 0.36±0.04 0.73±0.07 0.32±0.07
Simple DP3 [ze20243d] 1.00±0.00 0.59±0.01 0.45±0.03 0.79±0.07 0.26±0.04 0.63±0.07 0.22±0.07
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.58±0.06 0.53±0.12 0.80±0.06 0.38±0.09 0.70±0.06 0.28±0.05
BridgePolicy 1.00±0.00 0.84±0.07 0.58±0.10 0.88±0.05 0.44±0.07 0.76±0.07 0.32±0.07
Table 11: Success Rates of Individual Task on MetaWorld (Easy) benchmark.
MetaWorld (Easy)
[…]
button-
press
button-
press-
topdown
button-
press-
topdown-
wall
button-
press-
wall
coffee-
button
dial-
turn
door-
close
DP [chi2023diffusion] 0.93±0.01 0.98±0.01 0.96±0.03 0.93±0.03 0.99±0.01 0.63±0.10 1.00±0.00
DP3 [ze20243d] 1.00±0.00 1.00±0.00 0.99±0.02 0.99±0.01 1.00±0.00 0.66±0.01 1.00±0.00
Simple DP3 [ze20243d] 1.00±0.00 1.00±0.00 0.98±0.02 0.93±0.02 1.00±0.00 0.59±0.01 0.97±0.00
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.97±0.03 0.98±0.02 1.00±0.00 1.00±0.00 0.88±0.10 0.90±0.08
BridgePolicy 1.00±0.00 1.00±0.00 0.93±0.02 1.00±0.00 1.00±0.00 0.83±0.02 1.00±0.00
Table 12: Success Rates of Individual Task on MetaWorld (Easy) benchmark.
MetaWorld (Easy)
[…]
door-
lock
door-
open
lever-
pull
drawer-
close
reach
window-
close
window-
open
DP [chi2023diffusion] 0.75±0.08 0.98±0.03 0.98±0.03 1.00±0.00 0.93±0.03 0.99±0.01 1.00±0.00
DP3 [ze20243d] 0.98±0.02 0.86±0.01 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
Simple DP3 [ze20243d] 0.98±0.02 0.76±0.08 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.66±0.08 1.00±0.00 0.75±0.01 1.00±0.00 0.76±0.08 0.73±0.05
BridgePolicy 1.00±0.02 1.00±0.00 0.77±0.02 1.00±0.00 0.36±0.02 1.00±0.00 1.00±0.00
Table 13: Success Rates of Individual Task on MetaWorld (Easy) benchmark.
MetaWorld (Easy)
[…]
hand-
pull
hand-
pull-
side
door-
unlock
peg-
unplug-
side
drawer-
open
reach-
wall
plate-
slide-
back
DP [chi2023diffusion] 0.75±0.08 0.98±0.03 0.98±0.03 1.00±0.00 0.93±0.03 0.99±0.01 1.00±0.00
DP3 [ze20243d] 0.98±0.02 0.86±0.01 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
Simple DP3 [ze20243d] 0.98±0.02 0.76±0.08 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.66±0.08 1.00±0.00 0.75±0.01 1.00±0.00 0.76±0.08 0.73±0.05
BridgePolicy 1.00±0.00 0.92±0.01 1.00±0.00 0.93±0.04 1.00±0.00 0.80±0.05 1.00±0.00
Table 14: Success Rates of Individual Task on MetaWorld (Easy) and MetaWorld (Medium) benchmarks.
MetaWorld (Easy) MetaWorld (Medium)
[…]
plate-
slide-
back-side
plate-
slide
plate-
slide-side
basketball
bin-
picking
box-
close
hammer
DP [chi2023diffusion] 0.99±0.00 0.75±0.04 1.00±0.00 0.85±0.06 0.15±0.04 0.30±0.05 0.15±0.06
DP3 [ze20243d] 1.00±0.00 1.00±0.01 1.00±0.00 0.98±0.02 0.38±0.30 0.42±0.03 0.72±0.04
Simple DP3 [ze20243d] 0.99±0.00 1.00±0.01 1.00±0.00 0.95±0.04 0.28±0.16 0.38±0.05 0.62±0.09
FlowPolicy [zhang2025flowpolicy] 0.65±0.05 1.00±0.00 1.00±0.00 0.66±0.06 0.66±0.14 0.81±0.04 0.98±0.02
BridgePolicy 1.00±0.00 1.00±0.00 1.00±0.00 1.00±0.00 0.45±0.06 0.80±0.08 0.95±0.07
Table 15: Success Rates of Individual Task on MetaWorld (Medium) benchmark.
MetaWorld (Medium)
[…]
peg-
insert-
side
push-
wall
soccer
coffee-
pull
coffee-
push
sweep
sweep-
into
DP [chi2023diffusion] 0.34±0.07 0.20±0.03 0.14±0.04 0.34±0.07 0.67±0.04 0.18±0.08 0.10±0.04
DP3 [ze20243d] 0.67±0.07 0.51±0.08 0.18±0.03 0.85±0.03 0.94±0.03 0.96±0.03 0.17±0.05
Simple DP3 [ze20243d] 0.48±0.07 0.38±0.08 0.16±0.03 0.92±0.12 0.86±0.06 0.88±0.03 0.09±0.05
FlowPolicy [zhang2025flowpolicy] 0.70±0.09 0.63±0.08 0.33±0.05 0.96±0.02 0.61±0.06 0.70±0.04 0.31±0.02
BridgePolicy 0.77±0.06 0.87±0.03 0.31±0.05 0.97±0.01 0.87±0.03 0.97±0.02 0.40±0.03
Table 16: Success Rates of Individual Task on MetaWorld (Hard) benchmark.
MetaWorld (Hard)
[…]
pick-
out-of-
hole
pick-
place
assembly push
hand-
insert
DP [chi2023diffusion] 0.00±0.00 0.02±0.01 0.15±0.01 0.28±0.03 0.09±0.02
DP3 [ze20243d] 0.14±0.09 0.12±0.04 0.99±0.01 0.62±0.03 0.14±0.04
Simple DP3 [ze20243d] 0.08±0.06 0.12±0.06 0.79±0.01 0.32±0.03 0.12±0.05
FlowPolicy [zhang2025flowpolicy] 0.31±0.04 0.63±0.05 1.00±0.00 0.73±0.05 0.26±0.02
BridgePolicy 0.28±0.06 0.65±0.02 1.00±0.00 0.74±0.03 0.25±0.06
Table 17: Success Rates of Individual Task on MetaWorld (Very Hard) benchmark.
MetaWorld (Very Hard)
[…]
stick-
push
stick-
pull
shelf-
place
pick-
place-wall
disassemble
DP [chi2023diffusion] 0.63±0.03 0.11±0.02 0.11±0.03 0.05±0.01 0.43±0.07
DP3 [ze20243d] 0.97±0.04 0.27±0.08 0.19±0.10 0.35±0.08 0.75±0.04
Simple DP3 [ze20243d] 0.97±0.05 0.15±0.08 0.05±0.01 0.28±0.05 0.50±0.03
FlowPolicy [zhang2025flowpolicy] 1.00±0.00 0.56±0.03 0.40±0.03 0.95±0.03 0.88±0.03
BridgePolicy 1.00±0.00 0.82±0.03 0.25±0.03 0.96±0.05 0.92±0.03
E.4 More Sample Efficiency
We provide additional sample efficiency ablation results in this section shown in Figure 13.
Figure 13: Additional sample efficiency ablation results.